OpenStudy (readergirl12):

What is the value of b? 11+4b=8b-13 B=?? Help? :c

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[11+4b=8b-13\] right?

OpenStudy (misty1212):

1) subtract \(4b\) from both sides (now the b's are on the right) 2) add \(13\) to both sides, (now the numbers are on the left) 3) divide both sides by \(4\) to get \(b\)

OpenStudy (readergirl12):

4/4=1 and 8/4=2 .. Which one do I use?

OpenStudy (misty1212):

lets do the steps in order

OpenStudy (misty1212):

\[11+4b=8b-13\] subtract \(4b\) from both sides, what do you get?

OpenStudy (readergirl12):

11-4b? I'm sorry for being so dumb. I've never been good at math.. >.<

OpenStudy (misty1212):

it is ok we do it slow

OpenStudy (misty1212):

on the left you have \(4b\) and on the the right you have \(8b\)

OpenStudy (misty1212):

if you subtract \(4b\) from both sides, on the left the \(b's\) will be gone since \(4b-4b=0\)

OpenStudy (misty1212):

so it you start with \[11+4b=8b-13\] and subtract \(4b\) you get \[11+4b-4b=8b-4b-13\\ 11=4b-13\] all the b's are on the right side now

OpenStudy (misty1212):

so when i said "step one subtract \(4b\) from both sides" it turns \[11+4b=8b-13\] into \[11=4b-13\] now we are ready for step 2

OpenStudy (misty1212):

or i am ready are you?

OpenStudy (readergirl12):

Yep

OpenStudy (misty1212):

we got all the b's on the right, put the numbers on the left on the right is \(-13\) to get rid of that, add \(13\) to both sides what do you get?

OpenStudy (readergirl12):

24?

OpenStudy (misty1212):

yes !

OpenStudy (misty1212):

well actually you get \[24=4b\] not "24"

OpenStudy (misty1212):

now to get b alone, divide 24 by 4

OpenStudy (readergirl12):

6. c:

OpenStudy (misty1212):

that is it

OpenStudy (misty1212):

now it took three steps, lets write them again \[11+4b=8b-13\] subtract \(4b\)\[11=4b-13\] add 13\[24=4b\] divide by 4\[6=b\]

OpenStudy (readergirl12):

Mkay. Thanks.. :D